industry has 1000 competitive firms, each producing 50 tons of output. At the current market price of $10, half of the firms have a short-run supply curve with a slope of 1; the other half each have a short-run supply curve with slope 2. The short-run elasticity of market supply is

a) 3/10 b) 1/5 c) 2/5 d) none of the above


1
Expert's answer
2021-06-25T11:03:50-0400

Given Situation:

  • The short-run elasticity of market supply:

Firstly, the slope for the first half firm will be:

"slope \space 1=\frac{\Delta p}{\Delta Q}=1"

The slope for another half firm will be:

"slope \space 2=\frac{\Delta p}{\Delta Q}=2"

Now, at 50% the supply elasticity for firm 1 will be:

elasticity for first half "=\frac{p}{Q}\times\frac{\Delta p}{\Delta Q}"

"=\frac{10}{50}\times 1=\frac{1}{5}"

At 50% the supply elasticity for firm 2 will be:"=\frac{p}{Q}\times\frac{\Delta p}{\Delta Q}"

"=\frac{10}{50}\times 2=\frac{1}{10}"

Therefore,

The short-run elasticity of market supply will be:

short run elasticity"=\frac{1}{5}+\frac{1}{10}=\frac{2+1}{10}=\frac{3}{10}"

Thus, the correct option will be:

A) 3/10